### v8

@@ 1 Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Ma and Xue, 2024).
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2). This is partly due to the limited development of connected vehicle infrastructure and data platforms, which constrains the large-scale deployment of higher-level autonomous functions (Qu et al., 2025). In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments. Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions. Research related to vehicle trajectory analysis (Berghaus et al., 2024) and 3D scene understanding (Lim et al., 2021) has become essential for improving system performance and safety in real-world driving environments, providing valuable insights for the further enhancement of lane detection capabilities in these challenging scenarios.
@@
Earlier lane detection algorithms mainly relied on traditional computer vision methods, involving image preprocessing, feature extraction, and lane model fitting. Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering. Hou et al.(Hou et al., 2016) used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines. While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
With deep learning’s success in many traffic tasks(Guo et al., 2023; Yang et al., 2023a; Yuan et al., 2023; Zaman et al., 2025), it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization. Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018). However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation. Qin et al.(Qin et al., 2020) addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
@@
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS), supporting key functionalities such as lane departure warnings and autonomous navigation. While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents. Furthermore, the lack of specialized datasets tailored to foggy conditions has limited the development of effective deep learning-based solutions for these challenging scenarios.
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. In addition, we generate foggy versions of two widely-used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), creating FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@ 2 Related Work
In recent years, advances in deep learning have substantially improved lane detection performance. Building on these advances, current approaches for lane detection can be categorized into four main paradigms based on their modeling principles: segmentation-based, anchor-based, curve-based, and row-wise-based methods.
@@ 2.1 Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN(Pan et al., 2018) extends the traditional layer-by-layer convolution operation with a slice-by-slice convolution operation, which passes the information horizontally and vertically between the pixels within a layer, thus enhancing the continuity of lane line recognition. RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network that utilizes row and column multi-head self-attention to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
@@ 2.2 Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet(Tabelini et al., 2021b) introduces an end-to-end convolutional neural architecture that maps from images to polynomials representing each lane marker in the image, as well as domain lane polynomials and confidence scores for each lane. LSTR(Liu et al., 2021b) can also directly output the parameters of the lane shape model, has developed a Transformer-based network that takes into account non-local interactions, enabling it to capture the narrow structure of the lane lines and global context information. DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
@@ 2.3 Anchor-based Lane Detection
2.4 Row-wise-based Lane Detection
Row-wise methods predict lane positions for each image row. E2E-LMD(Yoo et al., 2020) first proposes the row-wise-based lane detection task and designes an end-to-end architecture that directly outputs the lane marker positions. UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet(Liu et al., 2021a) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
3 Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To address this gap, we have collected and constructed a foggy weather lane detection dataset called FoggyLane. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
@@ 3.1 FoggyLane
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube, Accessed: 2024), resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. We split the dataset into training and testing sets at a 2:1 ratio, ensuring balanced distribution across scene types.
Figure 4: The different road types of FoggyLane images examples.
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.5k 1.5k 1640 x 590
Comparing FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane (Pan et al., 2018) and CurveLanes (Xu et al., 2020) cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To address this gap, FoggyLane provides a comprehensive solution for foggy weather lane detection with instance-level annotations, available in both Tusimple (Tusimple, 2017) and CULane (Pan et al., 2018) formats for broader accessibility.
@@ 3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017). The CULane dataset (Pan et al., 2018) is a large-scale and challenging benchmark that covers diverse scenarios including highways, urban streets, and rural areas in Beijing. In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
@@
Figure 6: The FoggyCULane dataset examples.
To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) decomposes the image to obtain domain-invariant structure information. The Depth-specific Attention (DSA) predicts an attention map that helps to suppress depth-irrelevant structures by leveraging high-level semantic information extracted from the raw image. After passing through the DSA Module and STE Module, as shown in Fig. 5, the attention map and the structure map are multiplied to obtain a concise and comprehensive depth-specific structure map. The Depth Prediction (DP) is responsible for depth estimation.

### v9

@@
Abstract
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS). Existing lane detection algorithms generally perform well under favorable weather conditions. However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents. This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments. To address this, we introduce the FoggyLane dataset, captured in real-world foggy scenarios, and synthesize two additional datasets, FoggyCULane and FoggyTusimple, from existing popular lane detection datasets. Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple. Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
keywords:
ADAS, Intelligent vehicle, Foggy lane detection dataset, Foggy lane detection, Global attention, Edge enhancement
journal: Expert Systems With Applications
@@ Abstract
[1]organization=Guangdong Key Laboratory of Intelligent Transportation System, School of Intelligent Systems Engineering, Sun Yat-sen University, city=Guangzhou, postcode=510275, country=China
@@
[2]organization=College of Automotive Studies, Tongji University, city=Shanghai, postcode=201804, country=China
[…]
[3]organization=College of Transportation Engineering, Tongji University, city=Shanghai, postcode=201804, country=China
[…]
[4]organization=Department of Civil and Environmental Engineering, University of Alberta, city=Edmonton, state=Alberta, country=Canada
[…]
[5]organization=State Key Laboratory of Automotive Chassis Integration and Bionics, Jilin University, city=Changchun, postcode=130025, country=China
@@ 1 Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving(Chen et al., 2026).
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging(Sang and Norris, 2025). In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
@@
Although end-to-end autonomous driving technologies have made significant progress at Level 3 (L3) and higher, most current vehicles, particularly commercial and urban vehicles, are still at Level 0 (L0), Level 1 (L1), and Level 2 (L2). In these lower levels of autonomous driving, lane detection remains a crucial technology, especially in complex and adverse road environments. Most existing vehicles rely on Advanced Driver Assistance Systems (ADAS) to enhance safety in complex scenarios, especially in active warning, lane-keeping, and other functions. For example, Zhang and Zhu (2019) proposed an autonomous path tracking control method for intelligent electric vehicles based on lane detection and an optimal preview strategy, demonstrating that accurate lane recognition and trajectory tracking are fundamental to ensuring driving stability and safety at lower automation levels (Zhang and Zhu, 2019). Similarly, recent research highlights that coordinated lane-changing strategies in mixed traffic, such as the decentralized multi-vehicle decision-making framework developed by Alshakhs et al., are essential for improving both traffic efficiency and safety in environments where human-driven vehicles and connected automated vehicles coexist (Alshakhs et al., 2025). Therefore, this research not only focuses on end-to-end autonomous driving technologies but also aims to improve lane detection performance in complex environments, providing stronger safety support and driving assistance for L0, L1, and L2 vehicles in adverse conditions.
Earlier lane detection algorithms mainly relied on traditional computer vision methods. Son et al.(Son et al., 2015) first detected the vanishing point to select the ROI adaptively, then extracted lane markings based on color features and performed clustering. Hou et al.(Hou et al., 2016) used the Canny operator for edge detection and applied a progressive probabilistic Hough transform to detect lines. While these feature-based methods are efficient and suitable for real-time applications, they require manual parameter tuning and struggle in challenging conditions such as worn or occluded lane markings.
With deep learning’s success in many traffic tasks(Zhang et al., 2024b; Shunmuga Perumal et al., 2023; Preeti and Rana, 2024; Awan et al., 2025), it has become the mainstream approach for lane detection due to its strong feature extraction capabilities and better generalization. Deep learning offers higher robustness, adapting well to complex road environments and handling missing or occluded lane markings(Neven et al., 2018; Abualsaud et al., 2021; Yu et al., 2018; Pan et al., 2018). However, many networks rely on pixel-by-pixel classification, whereas lane detection does not require dense pixel segmentation. Qin et al.(Qin et al., 2020) addressed this by reframing lane detection as a row-based segmentation problem, reducing computational cost and improving detection speed.
@@
Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS. While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents. Furthermore, the lack of specialized datasets tailored to foggy conditions has limited the development of effective deep learning-based solutions for these challenging scenarios.
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. Additionally, we create foggy versions of two commonly used lane detection datasets, CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017), resulting in the FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@ 2 Related Work
In recent years, advances in deep learning have substantially improved lane detection performance. Building on these advances, current lane detection methods are divided into four main paradigms: segmentation-based, anchor-based, curve-based, and row-wise-based approaches, according to their modeling strategies.
@@ 2.1 Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet(Neven et al., 2018) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN(Pan et al., 2018) modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer. RESA(Zheng et al., 2021) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer(Zhao et al., 2024) presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
@@ 2.2 Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet(Tabelini et al., 2021b) proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane. Similarly, LSTR(Liu et al., 2021b) utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions. DBNet(Dai et al., 2024) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
@@ 2.3 Anchor-based Lane Detection
2.4 Row-wise-based Lane Detection
Row-wise methods predict lane positions for each image row. E2E-LMD(Yoo et al., 2020) introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly. UFLD(Qin et al., 2020) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet(Liu et al., 2021a) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
3 Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane (Pan et al., 2018) and Tusimple (Tusimple, 2017) datasets to test the generalizability of our methods.
Figure 4: The different road types of FoggyLane images examples.
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes (Aly, 2008) 1.2k None 640 x 480 Tusimple (Tusimple, 2017) 6.4k None 1280 x 720 LLAMAS (Behrendt and Soussan, 2019) 100k None 1276 x 717 Appolloscape (Huang et al., 2018) 110k None 3384 x 2710 BDD100k (Yu et al., 2020) 100k None 1280 x 720 CULane (Pan et al., 2018) 133k None 1640 x 590 CurveLanes (Xu et al., 2020) 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.5k 1640 x 590
@@ 3.1 FoggyLane
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos(YouTube, Accessed: 2024), resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
Com and paring FoggyLane with existing datasets, such as Caltech-Lanes(Aly, 2008) and Tusimple(Tusimple, 2017), which focus on single traffic scenarios, and LLAMAS(Behrendt and Soussan, 2019), which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape (Huang et al., 2018) and BDD100k (Yu et al., 2020) include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane (Pan et al., 2018) and CurveLanes (Xu et al., 2020) cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To fill this gap, FoggyLane offers a thorough solution for lane detection in foggy conditions with instance-level annotations.
@@
Figure 6: The FoggyCULane dataset examples. (a) is the original CULane dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
@@ 3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane(Pan et al., 2018) and TuSimple(Tusimple, 2017). The CULane dataset (Pan et al., 2018) serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing. In contrast, the TuSimple dataset (Tusimple, 2017) primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
@@
To calculate the corresponding $d(x)$, we use the S2R-DepthNet(Chen et al., 2021) that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details. The Depth-specific Attention (DSA) generates an attention map from the original image. Once processed through the DSA and STE Modules, as depicted in Fig. 5, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map. The Depth Prediction (DP) module is then responsible for estimating the depth.